# Robot-R1: Reinforcement Learning for Enhanced Embodied Reasoning in Robotics

### v1

@@ 5 Conclusion
We have proposed Robot-R1, a novel training framework that enhances the embodied reasoning capabilities of Large Vision-Language Models in robotic domains. Our approach trains models to predict the next state based on image observations and current states from expert demonstrations, using explicit reasoning processed optimized through reinforcement learning. Our extensive experiments demonstrate that models trained with Robot-R1 not only outperform Supervised Fine-Tuning (SFT) based approaches on embodied reasoning tasks but also exhibit consistent performance improvements across diverse embodied tasks while SFT methods often suffer from performance degradation.
@@ Acknowledgements
This paper was supported by RLWRD.
@@ B.1 Computing Cost
All experiments are conducted on a single node consisting of four A100 80GB GPUs. Training Robot-R1 task approximately 12 hours for 1 epoch using the 7.5K prompts.
@@ B.2 Prompt Template for Generating Training Dataset
In this section, we provide the prompt templates used to generate the training datasets. To generate the Multiple-Choice Question Answering (MCQA) dataset as denoted in Section 3.2, we utilize three different prompt templates: the waypoint prediction MCQA prompt template (Figure 5), the state prediction MCQA prompt template (Figure 6), and the primitive movement prediction MCQA prompt template (Figure 7). These templates take the current state, task instructions, and question about problem.
For generating the SFT dataset, we employ a prompt structure analogous to the MCQA format. Specifically, we use two different prompt templates for both Direct SFT and CoT SFT: the waypoint prediction SFT prompt template (Figure 8) and the current state prediction SFT prompt template (Figure 9). However, the SFT dataset, unlike the Robot-R1 dataset, includes not only prompts but also answers. The Direct SFT dataset is structured to generate immediate answers, while the CoT SFT dataset incorporates a reasoning path for the waypoint prediction task. The CoT SFT dataset is designed to perform high-level action planning, subtask reasoning, and primitive movement prediction sequentially before predicting the final answer. The high-level action planning and subtask reasoning components were defined using human-created scripts.
@@
Primitive Movement prediction MCQA # You are Franka Robot Assistant: Task Planning and Execution System
@@
Figure 7: Primitive movement prediction MCQA prompt template
@@ B.3 Benchmark Setup
  Task Type: "high_level"
@@
  Task Type: "subtask"
@@
  Task Type: "primitive"
@@
Evulation Rubric and LLM-as-Judge prompt Rubric 0 points: Meaning completely different from ground truth
@@ 9. Focus on whether the answer achieves the same functional outcome as the ground truth
Figure 12: Training metrics for Robot-R1 across two learning approaches: Open-End QA and MCQA. Open-End QA approach directly generats next waypoint values and use L1 distance for reward which is define between the predicted and ground-truth states. We conducted 2 epoch training for Open-End QA, and 1 epoch training for MCQA (note that all models reported in the paper are based on 1 epoch training). Due to differences in dataset sizes between the two tasks, the graphs are plotted based on training steps. (a) Shows accuracy metrics for the MCQA task (left axis) and 1 - L1 distance for Open-End QA (right axis). (b) Shows mean response length over training steps for both task types.
Appendix C Qualitative Results

### v2

@@ 5 Conclusion
We have proposed Robot-R1, a novel training framework that enhances the embodied reasoning capabilities of Large Vision-Language Models in robotic domains. Our approach trains models to predict the next state based on image observations and current states from expert demonstrations, using explicit reasoning process optimized through reinforcement learning. Our extensive experiments demonstrate that models trained with Robot-R1 not only outperform Supervised Fine-Tuning (SFT) based approaches on embodied reasoning tasks but also exhibit consistent performance improvements across diverse embodied tasks while SFT methods often suffer from performance degradation.
@@ Acknowledgements
This work was supported by Institute for Information & communications Technology Promotion(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST); No. RS-2024-00509279, Global AI Frontier Lab) and RLWRLD, Inc.
@@ B.1 Computing Cost
All experiments are conducted on a single node consisting of four A100 80GB GPUs. Training Robot-R1 task approximately 12 hours for 5 epoch using the 7.5K prompts.
@@ B.2 Prompt Template for Generating Training Dataset
In this section, we provide the prompt templates used to generate the training datasets. To generate the Multiple-Choice Question Answering (MCQA) dataset as denoted in Section 3.2, we utilize three different prompt templates: the waypoint prediction MCQA prompt template (Figure 4), the state prediction MCQA prompt template (Figure 5), and the movement prediction MCQA prompt template (Figure 6). These templates take the current state, task instructions, and question about problem.
For generating the SFT dataset, we employ a prompt structure analogous to the MCQA format. Specifically, we use two different prompt templates for both Direct SFT and CoT SFT: the waypoint prediction SFT prompt template (Figure 7) and the current state prediction SFT prompt template (Figure 8). However, the SFT dataset, unlike the Robot-R1 dataset, includes not only prompts but also answers. The Direct SFT dataset is structured to generate immediate answers, while the CoT SFT dataset incorporates a reasoning path for the waypoint prediction task. The CoT SFT dataset is designed to perform planning, high-level action reasoning, and movement prediction sequentially before predicting the final answer. The planning and high-level action reasoning components were defined using human-created scripts.
@@
Movement prediction MCQA # You are Franka Robot Assistant: Task Planning and Execution System
@@
Figure 6: Movement prediction MCQA prompt template
@@ B.3 Benchmark Setup
  Task Type: "planning"
@@
  Task Type: "high_level action"
@@
  Task Type: "movement"
@@
Evaluation Rubric and LLM-as-Judge prompt Rubric 0 points: Meaning completely different from ground truth
@@ 9. Focus on whether the answer achieves the same functional outcome as the ground truth
Appendix C Additional Evaluation and Analysis
In this section, we present additional evaluations and analyses to further validate the effectiveness of Robot-R1. For efficiency, all experiments in this section are trained for one epoch before evaluation.
Table 11: Robot-R1 Bridge Bench results
Model Planning High-level Action Movement Spatial
Gemini-1.5 Pro [48] 1.70 1.16 0.39 1.65
Qwen2.5-VL-7B-Ins [7] 1.20 0.89 0.46 1.53
w Robot-R1 (Ours) 1.38 1.10 0.45 1.56
Robot-R1 Bench on real robot. To further investigate whether the embodied reasoning capability learned by Robot-R1 can transfer to real-world robot settings, we extend the Robot-R1 Bench to the Robot-R1 Bridge Bench [50]. The problem construction and evaluation procedure remain identical to the original benchmark, but we use 100 randomly sampled images from real-robot demonstration videos in BridgeV2 [50]. Table 11 shows the evaluation results on the Robot-R1 Bridge Bench. Despite being trained only on RLBench simulation images, the model trained with Robot-R1 demonstrates notable performance gains, indicating that the high-level embodied reasoning ability acquired through Robot-R1 successfully transfers to real-robot environments. However, improvements in the embodied reasoning tailored for low-level action are relatively limited, which we attribute to differences in camera viewpoints and coordinate systems between RLBench and BridgeV2.
Table 12: VLABench LVLM Bench results
Model M&T Common- Sense Semantic Spatial Physical- Law Complex
Qwen2.5-VL-7B-Ins [7] 39.02 40.78 37.64 35.83 39.20 38.51
w Robot-R1 (Ours) 46.58 41.5 37.62 40.33 25.33 33.87
w CoT-SFT 0 0 0 0 0 0
Evaluation on VLABench. We further evaluate whether the model trained with Robot-R1 generalizes to other robot agent environments beyond RLBench, using the VLABench [66] evaluation pipeline under the CoT setup. As shown in Table 12, the performance improves on the Mesh & Texture (M&T) and Spatial tasks, while it decreases slightly on the Physical Law and Complex Reasoning tasks. We hypothesize that this performance drop stems from the fundamental gap between these task types and the abilities learned via Robot-R1. For instance, Physical Law tasks include questions such as “We have three objects with different densities. Choose the object with the smallest density,” while Complex Reasoning tasks include questions such as “Please rearrange these books in chronological order starting from the first published to the latest.” These tasks primarily require reasoning unrelated to embodied manipulation. In contrast, the M&T and Spatial tasks involve instructions such as “Add salt to the dish” or “Add the second condiment from the bottom to the dish,” which demand spatial understanding and movement prediction directly targeted by Robot-R1. These findings demonstrate that the embodied reasoning capabilities learned through Robot-R1 can be effectively transferred to new robotic environments beyond RLBench.
Table 13: Results of the cold start experiment on Robot-R1 Bench
Model Planning High-level Action Movement Spatial
Qwen2.5-VL-7B-Ins [7] 1.66 1.04 0.58 1.40
+ High Quality CoT-SFT 0.98 1.34 0.88 1.4
+ Robot-R1 (Ours) 1.22 1.44 0.88 1.43
Cold-start strategy. Training with high-quality SFT reasoning data followed by RL is a promising direction, as demonstrated in Guo et al. [17]. To evaluate whether Robot-R1 remains effective in such a cold-start scenario, we conducted experiments using GPT-4.1-mini [41] to generate high-quality CoT-SFT data and then applied Robot-R1 fine-tuning on top of the SFT trained model with high-quality CoT-SFT data. Specifically, the SFT data are generated by conditioning on human-annotated ground-truth planning and high-level action information, and prompting the model to produce CoT-style responses under the MCQA prompt used for RL. All SFT training settings followed those in the main experiment. As shown in Table 13, Robot-R1 demonstrates consistent performance improvements even on top of the SFT model, indicating the effectiveness on cold-start settings. ‘However, SFT-only training still shows decreased planning task performance and limited improvement in spatial tasks, implying that even with high-quality supervision, models remain vulnerable to forgetting and poor generalization.
Appendix D More Quantitative Results
In this section, we present quantitative metrics observed during the Robot-R1 training process. Specifically, we examine how the reward and response length change during training. Additionally, we include results for the Open-End approach trained Robot-R1 in Section 4.4. Open-End approach, unlike MCQA approach, directly generats next waypoint values and use L1 distance for reward which is define between the predicted and ground-truth states. Figure 11(a) illustrates how the MCQA accuracy and the accuracy reward obtained from 1-L1 distance change during the training process. These results show that GRPO training is being performed effectively. As shown in Figure 11(b), Robot-R1 also demonstrates that the response character length decreases throughout the training process. We hypothesize that this occurs as the reasoning process transitions from a summary format to a narrative format, during which redundant or unnecessary information is removed from the responses. Detailed analysis is provided in Section E.
@@ Appendix D More Quantitative Results
Figure 11: Training metrics for Robot-R1 across two learning approaches: Open-End QA and MCQA. We conducted 10 epoch training for Open-End QA and 5 epoch training for MCQA (note that all models reported in the paper are based on 5 epoch training). Due to differences in dataset sizes between the two tasks, the graphs are plotted based on training steps. (a) Shows accuracy metrics for the MCQA task (left axis) and 1 - L1 distance for Open-End QA (right axis). (b) Shows mean response length over training steps for both task types.
Appendix E Qualitative Results
@@
Table 14: VLA performance on LIBERO
Model Long Goal Objective Spatial Avg.
Qwen2.5-VL-7B-Ins [7] 44.0 61.6 83.4 86.8 69.0
w Robot-R1 (Ours) 45.8 73.8 80.4 87.0 71.8
Table 15: VLA performance on real robot
Model PnP 1 PnP 2 PnP 3 PnP 4 Avg.
Qwen2.5-VL-7B-Ins [7] 25 0 20.83 20.83 16.67
w Robot-R1 (Ours) 25 29.17 25 16.67 23.96
Appendix F Experiments on Vision-Language-Action Models
Finally, we evaluate whether training with Robot-R1 not only improves embodied reasoning but also enhances downstream robot control performance. We utilize the GR00T-N1.5 [16] repository for attaching a new diffusion policy head top on the LVLM backbone. Experiments are conducted in both the LIBERO simulation [35] environment and on a real robot.
LIBERO simulation. In LIBERO simulation, the environment is simulated using a Franka Panda Arm and includes four evaluation categories, such as, Spatial, Object, Goal, and Long. Each category consists of 10 tasks, and every task is evaluated over 50 rollouts. The action policy is trained using all given fine-tuning data with a batch size of 32 for 60k steps, following the original diffusion policy training setup while keeping all other hyperparameters identical to the GR00T-N1.5 repository. Table 14 presents the success rates in the four LIBERO categories. The model trained with Robot-R1 outperforms the baseline on average. In particular, the Goal category shows a remarkable performance improvement, which we attribute to its complex manipulation requirements that are closely aligned with the reasoning enhancements encouraged by Robot-R1. Although the Objective category shows a slight performance drop, we attribute this to the limited object diversity in RLBench during Robot-R1’s training stage.
Real robot experiment. In real robot experience, we utilize a Franka Research 3 arm to perform four different pick-and-place tasks, each involving four distinct objects: a teddy bear, box, cup, and sponge. Each task is evaluated 24 times, resulting in 96 total trials to calculate the success rate. For training, each object has 60 demonstration trajectories, leading to 240 training examples per task. For each task, we train a separate action diffusion policy using its corresponding demonstrations for 30k steps with a batch size of 32, while keeping all other hyperparameters consistent with the simulation setup. Table 15 shows that training the action diffusion policy with Robot-R1 improves the average success rate from 16.67% to 23.96%, confirming Robot-R1 effectiveness in enhancing real-world robot control performance.